i need to figure out the equation of 8x-10y=-80 and -2x-9y=18 then graph it to the y intercept and slope it and then figure out the intersection
Answer by mananth(16946)

8 x -10 y = -80 .............1
-2 x -9 y = 18 .............2

multiply (1)by -9
Multiply (2) by 10
-72 x + 90 y = 720
-20 x -90 y = 180
Add the two equations
-92 x = 900
/ -92
x = -9.78 ==> -88/9
plug value of x in (1)
8 x + -10 y = -80
-78.26 + -10 y = -80
-10 y = -80 + 78.26
-10 y = -1.74
y=1.74
Determine the intercep[t points of both the lines
equation1
when x= 0 y= 8 ( 0 , 8 )
when y = 0 x= -10 ( -10 , 0 )
Equation 2
when x= 0 y= -2 ( 0 , -2 )
when y = 0 x= -9 ( -9 , 0 )
y = 0.17 ==> 1/6



m.ananth@hotmail.ca